OpenStudy (anonymous):

limit as x approaches 0 tan^2x/x

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\tan^2x}{x}=\lim_{x\to0}\frac{\frac{\sin^2x}{\cos^2x}}{x}=\lim_{x\to0}\frac{\sin x}{x}\cdot\frac{\sin x}{\cos^2x}\]